# Statistical Inference with Limited Memory: A Survey

### v1

@@ 6.1 Randomized Binary Hypothesis testing
Hellman and Cover [84] addressed the problem of binary hypothesis testing within the class of randomized time-invariant finite-state machines, and they have found an exact characterization of the smallest attainable error probability for this problem. [^3] In the paper, the authors discuss the Bayesian setting, in which each hypothesis is assigned a prior probability. For consistency, we present the result for the minimax setting, which coincides with their result when the prior is equiprobable. Specifically, define the maximum and minimum likelihood ratios as
@@
Hellman and Cover showed that the error probability of any algorithm cannot be smaller than (23), and that, for any $\varepsilon>0$, there exists an $S$-states randomized estimation algorithm that achieves error probability of $\mathsf{P^{*}_{e}}(S)+\varepsilon$. As the optimal error probability decreases exponentially with $S$, we can define the asymptotics of the error exponent with regards to $S$ as
@@
For the lower bound, first assume that the Markov process induced by the finite-state algorithm is irreducible. There is no loss of generality in that assumption, as it is proved that the error probability of any $(S+1)$-state estimator that induces a reducible Markov process is lower bounded by $\mathsf{P^{*}_{e}}(S)$. For such processes, there is a unique stationary distribution induced by the input distribution. Denote by $\mu_{i}^{0}$ (resp. $\mu_{i}^{1}$) the stationary probability of state $i$ in the chain, under hypothesis $\mathcal{H}_{0}$ (resp. $\mathcal{H}_{1}$), and define the state likelihood ratios (SLR) $\lambda_{i}\triangleq\mu_{i}^{0}/\mu_{i}^{1}$. A low error probability intuitively implies that $\lambda_{i}\ll 1$ or $\lambda_{i}\gg 1$ for most $i\in[S]$, so to get a bound that holds for any machine we first need to show that the state likelihood ratio cannot grow too much. Specifically, assuming w.l.o.g. that the SLR’s are non-decreasing, it is shown that
@@
The upper bound construction can be understood by the following intuition: Assume that the hypotheses testing problem is dealing with the Bernoulli symmetric case, where the samples arrive either from a $\mathsf{Bern}(p)$ or a $\mathsf{Bern}(1-p)$ distribution. Consider a saturable counter machine, which is a machine with transition rule that moves up upon seeing $1$, moves down upon seeing 00, and either stays in place or moves back when it arrives at the extreme states (called saturated states). Furthermore, let the decision rule map each state to the hypothesis with the larger stationary distribution. A simple math exercise shows that the error probability we obtain decreases exponentially in the number of states, but at a rate that does not attains the lower bound described above.
The ingenious modification of Hellman and Cover is to move more probability mass to the saturated states, where an error is less likely. To achieve this, the algorithm only stays back from saturation when, in addition to the input sample pointing the other way, an external $\mathsf{Bern}(\delta)$ random coin falls on head. The parameter $\delta$ is typically very small, and its value depends on $p$. Extending it to the general binary hypothesis testing problem (not necessarily Bernoulli symmetric), the machine now moves up on high likelihood-ratio events, moves down on low likelihood-ratio events, and stays put otherwise. In the extreme state corresponding to hypothesis $\mathcal{H}_{0}$, the process moves back upon seeing a low likelihood-ratio event with probability $\delta$, chosen according to $\gamma$. In the extreme state corresponding to hypothesis $\mathcal{H}_{1}$, the process moves back upon seeing a high likelihood-ratio event with probability $k\cdot\delta$, where $k$ is chosen to account for the asymmetry between the hypotheses (see Figure 1). This defines a class of $\varepsilon$-optimal automata, i.e., for any $\varepsilon>0$, there exists an automaton in this class that achieves probability of error at most $\operatorname{\mathsf{P_{e}}}^{*}(S)+\varepsilon$. Thus, the optimal value can be approached arbitrarily closely.
@@ 6.2 Deterministic Binary Hypothesis testing
As stated previously, the binary hypothesis testing algorithm that achieves the Hellman-Cover lower bound is randomized. There have been some focus on the limited-memory hypothesis testing problem with restriction to deterministic algorithms. For simplicity, assume we observe either the $\mathsf{Bern}(p)$ distribution, under hypothesis $\mathcal{H}_{0}$, or the $\mathsf{Bern}(q)$ distribution, under hypothesis $\mathcal{H}_{1}$, for $10<q<p<$ 1. Since the optimal error probability decreases exponentially in $S$, we can define the error exponent as
@@
As deterministic machines are a special case of randomized machines, we clearly have $\mathsf{E}_{\mathsf{det}}(p,q)\leq\mathsf{E}(p,q)$. To demonstrate the important role randomization plays in approaching the optimal performance, Hellman and Cover showed in [96] that for any memory size $S<\infty$ and $\delta>0$ there exists problems such that any $S$-state deterministic machine has probability of error $\operatorname{\mathsf{P_{e}}}\geq\frac{1}{2}-\delta$, while the randomized machine from [84] has $\operatorname{\mathsf{P_{e}}}\leq\delta$. In [97] (see also [98]) it was shown that $\mathsf{E}_{\mathsf{det}}(p,q)$ is positive for all $p\neq q$, and for the symmetric setting, where $p=1-q$, Shubert et al. [99] have also derived a lower bound on $\mathsf{E}_{\mathsf{det}}(p,q)$. This implies that whenever $\gamma<\infty$, i.e., for any $10<p,q<$ 1, there exists some integer $1\leq C=C(p,q)<\infty$ such that $\operatorname{\mathsf{P^{*}_{e}}}_{\text{,det}}(S\cdot C)\leq\operatorname{\mathsf{P^{*}_{e}}}(S)$, for all $S$.
The problem has been largely abandoned at that point, until fairly recently, when Berg et al. [85] analyzed the error probability of optimal deterministic machines, and gave upper and lower bounds on $\mathsf{E}_{\mathsf{det}}(p,q)$ which coincide for some extreme values of the hypotheses. Specifically, they show that the optimal error probability of a deterministic machine has
@@
The lower bound on the error exponent is comprised of an algorithm that waits for events that very sharply distinguishes between the hypotheses, specifically, a long consecutive run of either zeros or ones, where the run length is chosen to optimize the error probability (see Figure 2). This $N$-state algorithm for testing between the hypotheses $\mathcal{H}_{0}=p$ and $\mathcal{H}_{1}=q$ is referred to as $\operatorname{\mathsf{RUNS}}(N,p,q)$, and it achieves error probability $\delta$ whenever $p=q+\varepsilon$ if the number of states is $N=O(1/\varepsilon\cdot\log(1/\delta))$.
@@
The upper bound is based on the insight that, while in randomized schemes the transition probabilities between states can be as small as desired, in deterministic machines the transition probabilities can only be as small as $\min\{p,1-p\}$ or $\min\{q,1-q\}$. More specifically, in order to derive a lower bound on $\operatorname{\mathsf{P^{*}_{e}}}_{\text{,det}}(S)$ for ergodic machines, it is first shown that the error probability of any algorithm is lower bounded by $\max_{i\in[S]}\min\{\mu_{i}^{p},\mu_{i}^{q}\}$. It is then left to show that there exits a state $i\in[S]$ for which both $\mu_{i}^{p}$ and $\mu_{i}^{q}$ are large. The proof follows the observation that, if state $i$ is connected to state $j$, we must have $\mu_{j}^{p}\geq\mu_{i}^{p}\cdot\min\{p,1-p\}$. Using the argument recursively, and applying it to $\mu_{i}^{q}$ as well, gives the bound.
For the non ergodic case, first consider the two absorbing states case, one for each competing hypothesis. It is shown that in any finite state machine, there must exist some state that is highly likely under both hypotheses and is close to both absorbing states, and this forces the error probability to be large as well. The general non ergodic case is proven by a superposition of the ergodic and two absorbing states cases. Though in general not tight, the lower bound demonstrates the gap between the error exponent for deterministic machines and that of randomized ones, which was derived in [84]. In particular, for certain values of $(p,q)$, it was shown that the restriction to deterministic machines may arbitrarily degrade the error exponent. Specifically, for any fixed $q<\frac{1}{2}$, we have $\lim_{p\rightarrow 1}\mathsf{E}_{\mathsf{det}}(p,q)=-\log q$ while $\lim_{p\rightarrow 1}\mathsf{E}(p,q)=\infty$, and, similarly, for any fixed $p>\frac{1}{2}$, we have $\lim_{q\rightarrow 0}\mathsf{E}_{\mathsf{det}}(p,q)=-\log(1-p)$ while $\lim_{q\rightarrow 0}\mathsf{E}(p,q)=\infty$.
@@ 6.3 Parity Learning
Following a conjecture of Steinhardt et al. [83], Raz proved in a seminal paper [86] that the problem of parity learning provides the following separation: Any algorithm for parity learning that uses less than $d^{2}/25$ bits of memory requires at least $2^{d/25}$ samples. To appreciate this result, assume that $\delta$ is some small constant. First note that for the parity learning problem we have $\mathsf{SC}(\delta)\asymp d$ and $\mathsf{MC}(\delta)\asymp d$ in bits. The sample-complexity follows since each observation gives at most $1$ bit of information on $\theta$, and $O(d)$ time can be achieved by using the natural approach of simply saving all the equations as if there are no memory limitations using $O(d^{2})$ memory bits, and then performing Gaussian elimination. The memory-complexity follows since we need at least $d$ bits to represent $\theta$, and we can achieve $O(d)$ by the rare events approach of waiting for all the singletons, which takes an exponential number of samples in $d$. The results of [86] can be translated as a sharp decay of $n^{*}(S,\delta)$ on the sample-memory curve, around $S\asymp d^{2}$, from exponential in $d$ to linear in $d$.
@@
Building on this result, Kol et al. [100] considered the problem of learning sparse parities. They have shown that the sparse parity learning problem, over the set of all vectors of Hamming weight exactly $\ell$, requires either a memory of size super-linear in $d$ or a number of samples super-polynomial in $d$, as long as $\ell\geq\omega(\log d/\log\log d)$. The work of Raz [101] expanded these previous results by showing that the same sharp drop-off of $n^{*}(S,\delta)$ in the sample memory curve is not only inherent to the parity learning problem, but is exhibited by a large class of problems. Specifically, let $\Theta,\mathcal{X}$ be two finite sets, and let $M:\Theta\times\mathcal{X}\rightarrow\{-1,1\}$ be a matrix. The learning problem is the following: An unknown element $\theta\in\Theta$ is chosen uniformly at random. A learner then tries to learn $\theta$ from a stream of samples, $(X_{1},Y_{1}),(X_{2},Y_{2}),\ldots$, where for every $i,X_{i}\in\mathcal{X}$ is chosen uniformly at random and $Y_{i}=M(X_{i},\theta)$. Let $\sigma_{M}$ be the largest singular value of $M$ and note that $\sigma_{M}\leq|\mathcal{X}|^{\frac{1}{2}}\cdot|\Theta|^{\frac{1}{2}}$, since the spectral norm is upper bounded by the Frobenius norm. The paper shows that if $\sigma_{M}\leq|\mathcal{X}|^{\frac{1}{2}}\cdot|\Theta|^{\frac{1}{2}-\varepsilon}$, then any learning algorithm for the corresponding learning problem requires either a memory of at least $\Omega((\varepsilon d)^{2})$ bits or at least $2^{\Omega(\varepsilon d)}$ samples, where $d=\log|\Theta|$.
@@
In a recent work, Garg et al. [102] leveraged the tools developed in [86] to give memory-sample lower bounds for learning parity with noise. In particular, they showed that when a stream of random linear equations over $\mathbb{F}_{2}$ is of the form $(X_{1},\theta^{T}X_{1}\oplus Z_{1}),\ldots,(X_{n},\theta^{T}X_{n}\oplus Z_{n})$ where $\{Z_{i}\}_{i=1}^{n}$ are distributed i.i.d $\mathsf{Bern}(\varepsilon)$ for some $20<\varepsilon<1/$ 2, then the drop-off of $n^{*}(S,\delta)$ in the sample-memory curve occurs around $S\asymp d^{2}/\varepsilon$.
@@
Assume that $\dim(E_{v})\geq k$ and let $Z_{0},\ldots,Z_{m}$ be the vertices on the computation path and $r_{i}=\dim(E_{v}\cap E_{Z_{i}})$. Note that $\dim(E_{j})\leq\dim(E_{i})+1$ implies $r_{i+1}\leq r_{i}+1$ for all $i$. Thus, if we reach a $v$ with $\dim(E_{v})\geq k$, then $r_{i+1}>r_{i}$ occurs at least $k$ times, so it is enough to bound $\Pr(r_{i+1}>r_{i})$. For this event to occur, the new sample $(x_{i+1},y_{i+1})$ must lay in the space spanned by $E_{v}$, hence it can be shown that $\Pr(r_{i+1}>r_{i})\leq 2^{-\Omega(d)}$. Then, by using the union bound over $S^{k}$ vertices, the probability to reach any $v$ with $\dim(E_{v})\geq k$ is less than $S^{k}\cdot(2^{-\Omega(d)})^{k}=2^{-\Omega(d^{2})}$.

### v2

@@ 6.1 Randomized Binary Hypothesis testing
Hellman and Cover [5] addressed the problem of binary hypothesis testing within the class of randomized time-invariant finite-state machines, and found an exact characterization of the smallest attainable error probability for this problem. [^4] In the paper, the authors discuss the Bayesian setting, in which each hypothesis is assigned a prior probability. For consistency, we present the result for the minimax setting, which coincides with their result when the prior is equiprobable. Specifically, define the maximum and minimum likelihood ratios as
@@
Hellman and Cover showed that the error probability of any algorithm cannot be smaller than (23), and that, for any $\varepsilon>0$, there exists an $S$-states randomized algorithm that achieves error probability of $\mathsf{P^{*}_{e}}(S)+\varepsilon$. As the optimal error probability decreases exponentially with $S$, we can define the asymptotics of the error exponent with regards to $S$ as
@@
For the lower bound, first assume that the Markov process induced by the finite-state algorithm is irreducible. There is no loss of generality in that assumption, as it is proved that the error probability of any $(S+1)$-state algorithm that induces a reducible Markov process is lower bounded by $\mathsf{P^{*}_{e}}(S)$. For such processes, there is a unique stationary distribution induced by the input distribution. Denote by $\mu_{i}^{0}$ (resp. $\mu_{i}^{1}$) the stationary probability of state $i$ in the chain, under hypothesis $\mathcal{H}_{0}$ (resp. $\mathcal{H}_{1}$), and define the state likelihood ratios (SLR) $\lambda_{i}\triangleq\mu_{i}^{0}/\mu_{i}^{1}$. A low error probability intuitively implies that $\lambda_{i}\ll 1$ or $\lambda_{i}\gg 1$ for most $i\in[S]$, so to get a bound that holds for any machine we first need to show that the state likelihood ratio cannot change too much. Specifically, assuming w.l.o.g. that the SLR’s are non-decreasing, it is shown that
@@
The upper bound construction can be understood by the following intuition: Assume that the hypothesis testing problem is dealing with the Bernoulli symmetric case, where the samples arrive either from a $\mathsf{Bern}(p)$ or a $\mathsf{Bern}(1-p)$ distribution. Consider a saturable counter machine, which is a machine with transition rule that moves up upon seeing $1$, moves down upon seeing 00, and either stays in place or moves back when it arrives at the extreme states (called saturated states). Furthermore, let the decision rule map each state to the hypothesis with the larger stationary distribution. A simple math exercise shows that the error probability we obtain decreases exponentially in the number of states, but at a rate that does not attains the lower bound described above.
The ingenious modification of Hellman and Cover is to move more probability mass to the saturated states, where an error is less likely. To achieve this, the algorithm only stays back from saturation when, in addition to the input sample pointing the other way, an external $\mathsf{Bern}(\delta)$ random coin falls on head. The parameter $\delta$ is typically very small, and its value depends on $p$. Extending it to the general binary hypothesis testing problem (not necessarily Bernoulli symmetric), the machine now moves up on high likelihood-ratio events, moves down on low likelihood-ratio events, and stays put otherwise. In the extreme state corresponding to hypothesis $\mathcal{H}_{0}$, the process moves back upon seeing a low likelihood-ratio event with probability $\delta$, chosen according to $\gamma$. In the extreme state corresponding to hypothesis $\mathcal{H}_{1}$, the process moves back upon seeing a high likelihood-ratio event with probability $k\cdot\delta$, where $k$ is chosen to account for the asymmetry between the hypotheses (see Figure 1). This defines a class of $\varepsilon$-optimal finite state machines, i.e., for any $\varepsilon>0$, there exists an machine in this class that achieves probability of error at most $\operatorname{\mathsf{P_{e}}}^{*}(S)+\varepsilon$. Thus, the optimal value can be approached arbitrarily closely.
@@ 6.2 Deterministic Binary Hypothesis testing
As stated previously, the binary hypothesis testing algorithm that achieves the Hellman-Cover lower bound is randomized. In the following, we restrict our discussion to deterministic algorithms. For simplicity, assume we observe either the $\mathsf{Bern}(p)$ distribution, under hypothesis $\mathcal{H}_{0}$, or the $\mathsf{Bern}(q)$ distribution, under hypothesis $\mathcal{H}_{1}$, for $10<q<p<$ 1. Since the optimal error probability decreases exponentially in $S$, we can define the error exponent as
@@
As deterministic machines are a special case of randomized machines, we clearly have $\mathsf{E}_{\mathsf{det}}(p,q)\leq\mathsf{E}(p,q)$. To demonstrate the important role randomization plays in approaching the optimal performance, Hellman and Cover showed in [101] that for any memory size $S<\infty$ and $\delta>0$ there exists problems such that any $S$-state deterministic machine has probability of error $\operatorname{\mathsf{P_{e}}}\geq\frac{1}{2}-\delta$, while the randomized machine from [5] has $\operatorname{\mathsf{P_{e}}}\leq\delta$. In [102] (see also [103]) it was shown that $\mathsf{E}_{\mathsf{det}}(p,q)$ is positive for all $p\neq q$, and for the symmetric setting, where $p=1-q$, Shubert et al. [104] also derived a lower bound on $\mathsf{E}_{\mathsf{det}}(p,q)$. This implies that whenever $\gamma<\infty$, i.e., for any $10<p,q<$ 1, there exists some integer $1\leq C=C(p,q)<\infty$ such that $\operatorname{\mathsf{P^{*}_{e}}}_{\text{,det}}(S\cdot C)\leq\operatorname{\mathsf{P^{*}_{e}}}(S)$, for all $S$.
The problem has been largely abandoned at that point, until fairly recently, when Berg et al. [90] analyzed the error probability of optimal deterministic machines, and gave upper and lower bounds on $\mathsf{E}_{\mathsf{det}}(p,q)$ which coincide for some extreme values of the hypotheses. Specifically, they showed that the optimal error probability of a deterministic machine has
@@
The lower bound on the error exponent is comprised of an algorithm that waits for events that very sharply distinguish between the hypotheses, specifically, a long consecutive run of either zeros or ones, where the run length is chosen to optimize the error probability (see Figure 2). This $N$-state algorithm for testing between the hypotheses $\mathcal{H}_{0}=p$ and $\mathcal{H}_{1}=q$ is referred to as $\operatorname{\mathsf{RUNS}}(N,p,q)$, and it achieves error probability $\delta$ whenever $p=q+\varepsilon$ if the number of states is $N=O(1/\varepsilon\cdot\log(1/\delta))$.
@@
The upper bound is based on the insight that, while in randomized schemes the transition probabilities between states can be as small as desired, in deterministic machines the transition probabilities can only be as small as $\min\{p,1-p\}$ or $\min\{q,1-q\}$. More specifically, in order to derive a lower bound on $\operatorname{\mathsf{P^{*}_{e}}}_{\text{,det}}(S)$ for ergodic machines, it is first shown that the error probability of any algorithm is lower bounded by $\max_{i\in[S]}\min\{\mu_{i}^{p},\mu_{i}^{q}\}$, where $\mu_{i}^{p}$ is the stationary distribution of state $i$ when the samples are drawn from the $\mathsf{Bern}(p)$ distribution. It is then left to show that there exits a state $i\in[S]$ for which both $\mu_{i}^{p}$ and $\mu_{i}^{q}$ are large. The proof follows from the observation that, if state $i$ is connected to state $j$, we must have $\mu_{j}^{p}\geq\mu_{i}^{p}\cdot\min\{p,1-p\}$. Using the argument recursively, and applying it to $\mu_{i}^{q}$ as well, gives the bound. The authors then consider the two absorbing states case, one for each competing hypothesis. It is shown that in any finite state machine, there must exist some state that is highly likely under both hypotheses and is close to both absorbing states, and this forces the error probability to be large as well. The general non ergodic case is proven by a superposition of the ergodic and two absorbing states cases.
Though in general not tight, the lower bound demonstrates the gap between the error exponent for deterministic machines and that of randomized ones, which was derived in [5]. In particular, for certain values of $(p,q)$, it is shown that the restriction to deterministic machines may arbitrarily degrade the error exponent. Specifically, for any fixed $q<\frac{1}{2}$, we have $\lim_{p\rightarrow 1}\mathsf{E}_{\mathsf{det}}(p,q)=-\log q$ while $\lim_{p\rightarrow 1}\mathsf{E}(p,q)=\infty$, and, similarly, for any fixed $p>\frac{1}{2}$, we have $\lim_{q\rightarrow 0}\mathsf{E}_{\mathsf{det}}(p,q)=-\log(1-p)$ while $\lim_{q\rightarrow 0}\mathsf{E}(p,q)=\infty$.
@@ 6.3 Parity Learning
Following a conjecture of Steinhardt et al. [89], Raz proved in a seminal paper [91] that the problem of parity learning provides the following separation: Any algorithm for parity learning that uses less than $d^{2}/25$ bits of memory requires at least $2^{d/25}$ samples. To appreciate this result, assume that $\delta$ is some small constant. First note that for the parity learning problem we have $\mathsf{SC}(\delta)\asymp d$ and $\mathsf{MC}(\delta)\asymp d$ in bits. The sample-complexity follows since each observation gives at most $1$ bit of information on $\theta$, and $O(d)$ time can be achieved by using the natural approach of simply saving all the equations as if there are no memory limitations using $O(d^{2})$ memory bits, and then performing Gaussian elimination. The memory-complexity follows since we need at least $d$ bits to represent $\theta$, and we can achieve $O(d)$ by the rare events approach of waiting for all the singletons, which takes an exponential number of samples in $d$. The results of [91] can be translated as a sharp decay of $n^{*}(S,\delta)$ on the sample-memory curve, around $S\asymp d^{2}$ bits, from exponential in $d$ to linear in $d$.
@@
Building on this result, Kol et al. [105] considered the problem of learning sparse parities. They showed that the sparse parity learning problem, over the set of all vectors of Hamming weight exactly $\ell$, requires either a memory of size super-linear in $d$ or a number of samples super-polynomial in $d$, as long as $\ell\geq\omega(\log d/\log\log d)$. The work of Raz [106] expanded these previous results by showing that the same sharp drop-off of $n^{*}(S,\delta)$ in the sample memory curve is not only inherent to the parity learning problem, but is exhibited by a large class of problems. Specifically, let $\Theta,\mathcal{X}$ be two finite sets, and let $M:\Theta\times\mathcal{X}\rightarrow\{-1,1\}$ be a matrix. The learning problem is the following: An unknown element $\theta\in\Theta$ is chosen uniformly at random. A learner then tries to learn $\theta$ from a stream of samples, $(X_{1},Y_{1}),(X_{2},Y_{2}),\ldots$, where for every $i,X_{i}\in\mathcal{X}$ is chosen uniformly at random and $Y_{i}=M(X_{i},\theta)$. Let $\sigma_{M}$ be the largest singular value of $M$ and note that $\sigma_{M}\leq|\mathcal{X}|^{\frac{1}{2}}\cdot|\Theta|^{\frac{1}{2}}$, since the spectral norm is upper bounded by the Frobenius norm. The paper shows that if $\sigma_{M}\leq|\mathcal{X}|^{\frac{1}{2}}\cdot|\Theta|^{\frac{1}{2}-\varepsilon}$, then any learning algorithm for the corresponding learning problem requires either a memory of at least $\Omega((\varepsilon d)^{2})$ bits or at least $2^{\Omega(\varepsilon d)}$ samples, where $d=\log|\Theta|$.
@@
In a recent work, Garg et al. [107] leveraged the tools developed in [91] to give memory-sample lower bounds for learning parity with noise. In particular, they showed that when a stream of random linear equations over $\mathbb{F}_{2}$ is of the form $(X_{1},\theta^{T}X_{1}\oplus Z_{1}),\ldots,(X_{n},\theta^{T}X_{n}\oplus Z_{n})$ where $\{Z_{i}\}_{i=1}^{n}$ are distributed i.i.d $\mathsf{Bern}(\varepsilon)$ for some $20<\varepsilon<1/$ 2, then the drop-off of $n^{*}(S,\delta)$ in the sample-memory curve occurs around $S\asymp d^{2}/\varepsilon$ bits.
@@
Assume that $\dim(E_{v})\geq k$ and let $Z_{0},\ldots,Z_{m}$ be the vertices on the computation path and $r_{i}=\dim(E_{v}\cap E_{Z_{i}})$. Note that $\dim(E_{Z_{i+1}})\leq\dim(E_{Z_{i}})+1$ implies $r_{i+1}\leq r_{i}+1$ for all $i$. Thus, if we reach a $v$ with $\dim(E_{v})\geq k$, then $r_{i+1}>r_{i}$ occurs at least $k$ times, so it is enough to bound $\Pr(r_{i+1}>r_{i})$. For this event to occur, the new sample $(x_{i+1},y_{i+1})$ must lay in the space spanned by $E_{v}$, hence it can be shown that $\Pr(r_{i+1}>r_{i})\leq 2^{-\Omega(d)}$. Then, by using the union bound over $S^{k}$ vertices, the probability to reach any $v$ with $\dim(E_{v})\geq k$ is less than $S^{k}\cdot(2^{-\Omega(d)})^{k}=2^{-\Omega(d^{2})}$.